OpenStudy (anonymous):

The average of 11 numbers is 21 . The average of seven of those numbers is 25. What is the average of the remaining four numbers?

OpenStudy (mathmagician):

\[(7*25+x*4)/11=21\] \[4x=231-175=56\] x=14, so the average of the remaining 4 numbers is 14

OpenStudy (callisto):

total of the 11 number = 11x21 = 231 total of the seven number = 25*7 = 175 total of the 4 number = 231-175 = 56 average required = 56/4 = 14

OpenStudy (anonymous):

x1 + x2 + x3 ......+x11 = 21*11 x1 + x2 ...... +x7 = 25*7 Therefore x8 + x9 + x10 + x11 = 21*11 - 25*7 Average = (21*11 - 25*7 )/ 4 =56/4 = 14